6. Consider the following utility function: 𝒖 = πŸ’πŸŽπŸŽ (𝟏 + 𝟏)βˆ’πŸ. π’™π’š

a. Is this utility function homogenous? Briefly explain. [3]

b. Use implicit differentiation to find the 𝑀𝑅𝐢𝑆. [2]

c. Write down an expression for the indifference curve if 𝒖 = 𝟐𝟎. [2]


1
Expert's answer
2021-06-08T18:45:41-0400

(a) The function is not homogeneous

as "u=400[\frac{1}{x}+\frac{1}{y}]^{-1}"


"=\frac{400}{(y)^{-1}}[1-\frac{y}{x}+\frac{y^3}{x^2}-\frac{y^3}{x^3}+......]"


"=40y[1-\frac{y}{x}+\frac{y}{x}^2-\frac{y}{x}^3+......]"

not homogeneous.


(b)"mu=\frac{\delta u}{\delta x}(x,y)=400(-1)[(\frac{1}{x}+\frac{1}{y})]^{-2}"

"=(\frac{-1}{x^2})"

similarly

"mu_2=\frac{\delta u}{\delta x}(x,y)=400(-1)[(\frac{1}{x}+\frac{1}{y})]^{-2}[\frac{-1}{y^2}]"


so "MRS=\frac{\delta u}{\delta x}=\frac{\delta u/\delta x}{\delta u/\delta y}"


"=\frac{-Mu_1}{Mu_2}=\frac{-y^2}{x^2}"


"MRS=\frac{y^2}{ x^2}"


(c)given that u=20

so, "u=20=400[(\frac{1}{x}+\frac{1}{y})]^{-1}"

the required indifference curve is

"\frac{1}{x}+\frac{1}{y}=20"